A golfer hits a ball that reaches a height of 24 metres after 2 seconds. If the ball stays in the air for a total of 6 seconds, what was its maximum height?
Answer by khwang(438)

Let Vo be the initial vertical velocity of the ball.

Since h = vot - 1/2 gt^2,
and we know h = 24 m when t = 2 sec, we have
24 = 2 Vo - 2*9.8 ,
2 Vo = 24 + 19.6 = 43.6,
So Vo = 21.8 m/s

By the formula V^2 = Vo^2 - 2g h and so h is maximum when V =0,
so h(max) = Vo^2 / 2g = (21.8)^2 / 19.6 = 24.25 meters
Answwer: Maximum height is 24.25 meters